### 08 Apr 2026 v2

Abstract
Primitive-based methods such as 3D Gaussian Splatting have recently become the state-of-the-art for novel-view synthesis and related reconstruction tasks. Compared to neural fields, these representations are more flexible, adaptive, and scale better to large scenes. However, the limited expressivity of individual primitives makes modeling high-frequency detail challenging. We introduce Neural Harmonic Textures, a neural representation approach that anchors latent feature vectors on a virtual scaffold surrounding each primitive. These features are interpolated within the primitive at ray intersection points. Inspired by Fourier analysis, we apply periodic activations to the interpolated features, turning alpha blending into a weighted sum of harmonic components. The resulting signal is then decoded in a single deferred pass using a small neural network, significantly reducing computational cost. Neural Harmonic Textures yield state-of-the-art results in real-time novel view synthesis while bridging the gap between primitive- and neural-field-based reconstruction. Our method integrates seamlessly into existing primitive-based pipelines such as 3DGUT, Triangle Splatting, and 2DGS. We further demonstrate its generality with applications to 2D image fitting and semantic reconstruction.
Figure 1: Neural Harmonic Textures for novel view synthesis. We attach learnable feature vectors (right) to the virtual vertices of bounding tetrahedra encapsulating each primitive (center). After harmonic encoding and accumulation along the ray, a small neural network decodes the resulting signal into RGB color in a deferred manner (left). Source code and further results are available at https://research.nvidia.com/labs/sil/projects/neural-harmonic-textures/.
1 Introduction
Since the introduction of 3D Gaussian Splatting [kerbl20233Dgaussians], (Lagrangian) primitive-based approaches have largely displaced (Eulerian) continuous neural radiance fields [sitzmann20, mueller2022instant, takikawa2023compact] for novel view synthesis [kerbl20233Dgaussians, trianglesplatting].
This shift is driven not only by significantly improved rendering speed, but also by the structural advantages of explicit representations. Primitive-based methods naturally adapt to scene detail, scale more gracefully, and readily support motion, deformation, and editing [4dgaussians, chen2024gaussianeditor]. Furthermore, they align well with feed-forward reconstruction pipelines [gslrm2024, liang2024btimer] and closely resemble widely adopted point-map representations [VGGT, wang2025pi].
Despite these advantages, the limited expressive power of individual primitives remains a fundamental bottleneck. Geometry and appearance are tightly coupled within each primitive, forcing high-frequency spatial detail to be represented by increasing the number of primitives, which directly increases memory consumption and hampers rendering speed. Directional appearance modeling is constrained in a similar manner. View-dependent effects are typically represented using low-order Spherical Harmonics, which are spectrally band-limited and poorly suited for high-frequency phenomena such as sharp specular highlights. Alternative angular parametric functions, such as spherical Gaussians, spherical Beta kernels [betasplatting], and spherical Voronoi [sphericalVoronoi] increase directional bandwidth, but leave the fundamental coupling between spatial support and appearance modeling unchanged. In contrast, neural radiance fields achieve high local representational capacity by combining positional encodings such as Fourier features [tancik2020fourier, mildenhall2021nerf] or multi-resolution hash grids [mueller2022instant] with a learned multi-layer perceptron (MLP) decoding. Recent works [neuralshells, radiancemesh] attempted to bridge both paradigms by employing primitives as acceleration structures for querying a global neural field. However, these hybrid approaches inherit key limitations of global neural representations, particularly with respect to motion, deformation, editability and scalability.
In this work, we introduce Neural Harmonic Textures, a novel primitive-based neural representation that increases the expressive power of individual primitives while retaining the advantages of Lagrangian representations. To this end, we reinterpret primitives as both geometric carriers and local positional encodings. Specifically, we attach learnable feature vectors to a virtual scaffold encapsulating each primitive (e.g., Gaussians or triangles) and interpolate them locally at ray–primitive intersection points. Compared to globally anchored feature representations, our formulation moves with the primitives and therefore naturally supports operations such as motion, deformation, and scene editing.
In prior work, such interpolated features are typically processed by a lightweight MLP to decode RGB values, followed by alpha compositing across intersected primitives to produce the final color, requiring numerous neural evaluations. Instead, we draw inspiration from the Fourier transform, where a complex signal is expressed as a sum of harmonic components (i.e., periodic functions with different amplitudes). We therefore apply periodic functions (sine and cosine) to the interpolated feature vectors before compositing them along the ray. Under this formulation, the activated features act as frequency components, while the primitive opacity modulates their amplitude (Footnote˜2). The resulting implicit signal in image space is sufficiently rich to be directly decoded using a lightweight MLP in a single evaluation per pixel, in a deferred-shading manner.
We validate Neural Harmonic Textures on standard novel view synthesis benchmarks, achieving state-of-the-art performance among both real-time and offline methods (Sec.˜5.1). Our proposed formulation is agnostic to the underlying primitive type: triangles, 2D or 3D Gaussians, tetrahedra, and other elements can be used interchangeably, enabling seamless integration into existing primitive-based pipelines while significantly increasing their representational capacity (Sec.˜5.2). Finally, by abstracting the encoded signal, our approach naturally extends to higher-dimensional signals. Joint modeling of color and semantic features becomes possible within a unified local representation, allowing us to exploit both spatial and cross-modal correlations for improved efficiency and compactness (Sec.˜5.2.2).
2 Related Work
2.1 Neural Fields
Neural fields have become a foundational model for representing complex, multi-scale signals across a wide range of domains [xie2022neural], including audio [luo2022learning], images [acorn, saragadam2023wire], 3D geometry [park2019deepsdf, wang2021neus, takikawa2021nglod, li2023neuralangelo], radiance fields [mildenhall2021nerf], scientific data [wu2022instant, wu2024distributed, gadirov2025hyperflint], and volumetric video[pumarola2021d, fridovich2023k]. These methods primarily differ in how they encode their inputs and parameterize their outputs. In most formulations, latent features are stored in a compressed representation and subsequently decoded by a neural network. This storage can be implicit within the network weights [mildenhall2021nerf, gaussianscalespaces, Lombardi21, neuralVolumes, kilonerf, sitzmann2020implicit, tancik2020fourier], or explicit through dedicated spatial data structures. Explicit representations typically offer improved scalability and performance, though often at increased memory cost. Common examples include voxel grids [acorn], triplanes [chen2022tensorf, fridovich2023k], point clouds [govindarajan2024lagrangian], hash grids [mueller2022instant], and meshes [thies2019]. In addition, explicit spatial structures are frequently used as acceleration mechanisms to skip empty regions and efficiently identify relevant query locations [mueller2022instant, neuralshells, neuralsparsevoxelfields]. More recently, neural fields have been combined with primitive-based methods by turning the primitives into tiny neural fields with closed-form antiderivatives, enhancing their individual modelling power [zhou2025splat].
Beyond the choice of spatial representation, prior work has shown that positional encodings play a crucial role in enabling neural networks to represent high-frequency signals in low-dimensional spatial domains [tancik2020fourier, mildenhall2021nerf, mueller2022instant, sitzmann2020implicit]. In contrast to globally anchored spatial structures such as hierarchical voxel grids (hashgrid encodings), which follow an Eulerian formulation, our approach is Lagrangian. We anchor latent feature vectors relative to explicit particles, enabling adaptive spatial resolution and naturally supporting explicit motion, deformation, and scene editing.
2.2 Primitive-based Representations for Neural Reconstruction
Primitive-based methods have recently gained significant traction for 3D reconstruction and novel view synthesis, largely driven by the success of 3D Gaussian Splatting [kerbl20233Dgaussians]. Numerous extensions have since explored variants of Gaussian primitives [4dgaussians, betasplatting, 3dgut, 3dgrt, condor2025, condor2026, kulhanek2025lodge, 2dgs] as well as alternative representations, including Voronoi cells [radiantfoam], triangle soups [trianglesplatting], and connected meshes [radiancemesh]. While offering high rendering performance, editability, and interpretability, these approaches typically exhibit poor memory scaling when fitting high-frequency detail due to their joint modeling of geometry and appearance.
Neural field methods, in contrast, aggregate information across the entire scene, enabling compact representations. Concurrent work [neuralshells, radiancevolumetricmeshes] attempts to bridge this gap by using geometric primitives as acceleration structures for neural field evaluation. However, these approaches lose some of the advantages of purely primitive-based representations, as the learned features remain globally anchored (e.g. through hashgrids) and do not follow the primitives under motion or deformation. Furthermore, they inherit the limitations of Eulerian encodings, namely poor scaling in large scenes and high-frequency detail modelling. In contrast, we propose that geometric primitives themselves, when appropriately parameterized, can serve as an effective positional encoding mechanism, and enable more efficient signal decoding (i.e. deferred shading).
2.3 Deferred Shading and Texturing in Radiance fields
Alternatives to direct color evaluation in novel view synthesis have been explored to improve expressivity, rendering quality, and efficiency. Early work introduced image-based neural deferred shading conditioned on 3D meshes [thies2019]. Later approaches extended this idea to neural radiance fields by baking appearance features into extracted geometry for real-time rendering [bakedsdf, chen2022mobilenerf, hedman2021snerg], removing the necessity to run neural networks during runtime. However, these methods rely on multi-stage pipelines that first optimize a global field, then extract geometry and bake appearance features. In contrast, our method jointly optimizes primitives and neural appearance from scratch, preserving volumetric flexibility and eliminating the need for a separate baking stage, while retaining high quality and speed.
Deferred shading has also been explored in the context of radiance field rendering using neural networks as deferred shaders. In Han et al. [han2023], global feature grids are queried, with features accumulated and decoded once in a deferred manner. Similarly, point-based neural renderers have recently mixed local neural decoding with deferred shading. Hahlbohm et al. [hahlbohm25] uses point-clouds as query points for a global hash-grid encoded feature field, decoded locally using a shallow neural network into implicit features. These are splatted to the camera plane and finally rendered through a convolutional neural network. However, the necessity of large-bandwidth neural networks to uplift the poor scalability and detail afforded by the spatial encoding made them slow and difficult to train; the former requiring guiding networks with regular NeRF-style rendering to steer the optimization, while the latter required expensive perceptual-loss supervision.
To improve expressivity without global feature grids, several primitive-based methods introduce per-primitive textures [bbsplat, xu2024texture, texturedgaussians, huang2024textured, wurster2024gabor, zhou20253dgabsplat]. However, highly expressive primitives can complicate optimization and degrade reconstruction quality. Instead, we decouple geometry and appearance by pairing local per-particle features with a shared lightweight neural decoder, in turn enabling the reduction of queries to the neural field to a single deferred pass. This allows particle density to more closely follow geometric complexity rather than being artificially increased to reproduce high-frequency appearance.
3 Preliminaries
We review primitive-based radiance fields, primitive-based volume rendering, and positional encodings, focusing on the aspects most relevant to our method.
Primitive-based Radiance Fields.
Primitive-based radiance fields represent complex scenes using an unstructured collection of 2D or 3D geometric primitives, such as anisotropic Gaussians [kerbl20233Dgaussians], oriented disks [2dgs], or triangles [trianglesplatting]. In Sec.˜4, we present our formulation on the example of 3D Gaussian primitives [kerbl20233Dgaussians, 3dgut] whose spatial response function
$\rho(\mathbf{x})=\exp\!\left(-\frac{1}{2}(\mathbf{x}-\boldsymbol{\mu})^{T}\boldsymbol{\Sigma}^{-1}(\mathbf{x}-\boldsymbol{\mu})\right),$ (1)
is defined by the particle center $\boldsymbol{\mu}\in\mathbb{R}^{3}$ and covariance matrix $\boldsymbol{\Sigma}\in\mathbb{R}^{3\times 3}$. To ensure positive semi-definiteness during optimization, the covariance is parameterized as $\boldsymbol{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T}$, where $\mathbf{R}\in SO(3)$ represents rotation and $\mathbf{S}\in\mathbb{R}^{3\times 3}$ encodes scaling. In practice, these parameters are stored as a quaternion $\mathbf{q}\in\mathbb{R}^{4}$ and a scale vector $\mathbf{s}\in\mathbb{R}^{3}$. Each primitive additionally carries an opacity parameter $\sigma\in\mathbb{R}$ and a view-dependent radiance function $\phi(\mathbf{d})$, commonly represented using spherical harmonics. Note that in this case, $\phi(\mathbf{d})$ depends only on the ray direction and not on the spatial position of the intersection point.
Primitive-based Volume Rendering.
Given a camera ray $\mathbf{r}(\tau)=\mathbf{o}+\tau\mathbf{d}$ with origin $\mathbf{o}\in\mathbb{R}^{3}$ and direction $\mathbf{d}\in\mathbb{R}^{3}$, the rendered ray color $\mathbf{c}\in\mathbb{R}^{3}$ is obtained via volume rendering over the primitives $P_{c}$ intersecting the ray:
$\mathbf{c}(\mathbf{o},\mathbf{d})=\sum_{i\in P_{c}}\phi(\mathbf{d})\cdot T_{i}\cdot\alpha_{i},\quad T_{i}=\prod_{j=1}^{i-1}1-\alpha_{j},$ (2)
where the opacity $\alpha_{i}$ is defined as $\alpha_{i}=\sigma_{i}\,\rho_{i}(\mathbf{o}+\tau\mathbf{d})$ and $T_{i}$ is the transmittance.
Positional Encoding:
Neural networks exhibit a natural spectral bias that limits their ability to learn high-frequency functions in low-dimensional domains [tancik2020fourier].
Positional encodings mitigate this limitation by mapping low-dimensional spatial and directional coordinates into higher-dimensional representations before feeding them into the network [tancik2020fourier]. While early approaches relied on fixed frequency encodings [harris2010digital, muller2019neural], modern methods often use parametric encodings that store learnable feature vectors in auxiliary spatial data structures, such as grids or trees [mueller2022instant, takikawa2021nglod]. Query coordinates retrieve and interpolate these features, effectively shifting much of the representational burden from the neural network to the encoding itself. This design trades increased memory usage for reduced computation: only a small subset of encoding parameters is updated per sample, allowing the neural network to remain compact and efficient while significantly accelerating convergence and maintaining high reconstruction quality.
Figure 2: Neural Harmonic Textures applied to novel-view synthesis. We virtually attach feature vectors $\mathbf{f}_{i}$ to the vertices of tetrahedra inscribing the Gaussian primitives [^2] Gaussian particles are bounded by ellipsoids in world space, which become spheres after whitening. Our virtual bounding tetrahedra are defined in this canonical space. . Following 3DGUT [3dgut], we evaluate the point along the ray where the projected Gaussian has maximum response. We barycentrically interpolate vertex features at that point, and encode them with sine and cosine functions into different channels. These are then alpha blended along the rest of the ray, until the resulting sum of harmonics is decoded by a shallow MLP in a single image-space pass.
4 Method
We present the core components of our approach using 3D Gaussian primitives in the context of novel view synthesis, following the formulation of 3DGUT [3dgut]. We begin by introducing our primitive-bound feature embedding in Sec.˜4.1. Next, Sec.˜4.2 presents harmonic texturing, which increases per-primitive expressive power while preserving the locality and explicit form of the representation. Finally, Sec.˜4.3 describes a neural deferred decoding scheme that enables efficient signal reconstruction in a single image-space pass. Generalization to other primitive types and additional applications will be discussed in Sec.˜5.2. Footnote˜2 provides a schematic overview of our method.
(a) Latent vectors
(b) Interpolation
(c) Harmonic decomposition
Figure 3: Illustrating our method in 2D. Each primitive is bounded by an ellipsoid in world space, which becomes a sphere in whitened canonical space (a). Considering a virtual bounding tetrahedron in this canonical space, we attach one N-dimensional feature vector $\mathbf{f}^{j}$ to each vertex. The primitive’s contribution is evaluated at the point of maximum response $\mathbf{p}^{*}$ of the projected Gaussian along the intersecting ray (b). The feature vectors are barycentrically interpolated at $\mathbf{p}^{*}$ and encoded with sine and cosine periodic functions (c).
4.1 Primitive-bound feature embedding
Spatial structures such as triplanes [duckworth2023smerf], voxel grids [SunSC22], and multi-resolution hash grids [mueller2022instant] are commonly used in neural fields to hold latent feature vectors that embed query positions into a higher-dimensional space before decoding the signal. Despite their effectiveness, these representations rely on globally defined regular grids, which limits their scalability to large scenes or high-frequency detail. Moreover, because these structures are fixed in space, they struggle to represent scenes undergoing motion, deformation, or editing.
Instead, we exploit the Lagrangian nature and adaptivity of primitive-based representations by anchoring latent features to a virtual scaffold that encapsulates each primitive. Specifically, consider the isosurface ellipsoid defined by an anisotropic 3D Gaussian (Eq.˜1) with bounded support, which becomes a sphere in canonical space after whitening. We define the virtual tetrahedron bounding this sphere and assign a feature vector $\mathbf{f}^{j}\in\mathbb{R}^{N_{f}}$ to each of its four vertices ($j\in\{0..3\}$) (Fig.˜3a). For each primitive intersected by a ray, we determine the point $\mathbf{p}^{*}$ corresponding to the maximum Gaussian response along the ray [3dgut]. The feature vector at this location, $\mathbf{f}$, is obtained via barycentric interpolation of the vertex features $\mathbf{f}^{j}$. A detailed derivation is provided in the supplementary material.
Compared to globally anchored feature representations, our formulation moves with the primitives and therefore naturally supports operations such as motion, deformation, or deletion during scene editing.
4.2 Harmonic Texturing
In previous work, feature vectors are typically passed through a lightweight neural network to decode the primitive appearance prior to volume rendering, requiring dozens of MLP evaluations per ray to produce the final pixel color [mueller2022instant, nestsplatting, radiancemesh]. In 3DGS, color decoding overhead (without neural networks) is reduced by approximating view-dependent color emission once per primitive, thereby reducing the complexity from the number of ray–primitive intersections to the number of primitives. However, this approximation assumes that the signal does not vary spatially within each primitive, which makes it unsuitable for our representation.
An alternative strategy is to blend the features along each ray and decode the signal in image space, in the spirit of deferred shading [hedman2021snerg]. Inspired by the Fourier transform, where a complex signal is expressed as a sum of harmonic components (i.e., periodic functions of different amplitudes), we instead encode the interpolated features using periodic functions before blending them along the ray. This transforms the features into frequencies and amplitudes, yielding a harmonic decomposition of the signal, which we term Harmonic Textures (Fig.˜3c).
Figure 4: Harmonic textures.
In this formulation, the interpolation function effectively acts as a frequency modulator: large differences between vertex features within a primitive produce rapidly oscillating, spatially varying textures. Each primitive additionally has a kernel-weighted opacity, which acts as the harmonic amplitude. This behavior is illustrated in the inset Fig.˜4 (without the amplitude weighting, to better visualize the harmonics).
4.3 Neural Deferred Shading
The rich high-dimensional signal yielded by the volume rendering of Harmonic Textures enables a more efficient decoding strategy. We concatenate the accumulated harmonics with the ray direction $\mathbf{d}$ encoded using second-degree spherical harmonics $\mathrm{SH}_{2}(\mathbf{d})\in\mathbb{R}^{9}$ as done in [mueller2022instant], and decode the final pixel color $\mathbf{c}$ using a shallow MLP, without further positional encoding. The resulting rendering equation is
$\mathbf{c}=\mathrm{MLP}_{\theta}\left(\sum_{i\in\mathcal{G}}\alpha_{i}\,T_{i}\begin{bmatrix}\sin(\mathbf{f}_{i})\\ \cos(\mathbf{f}_{i})\end{bmatrix},\;k\cdot\mathrm{SH}_{2}(\mathbf{d})\right),$ (3)
where $\mathbf{f}_{i}=\mathrm{interpolate}\!\left(\left[\mathbf{f}_{i}^{0},\mathbf{f}_{i}^{1},\mathbf{f}_{i}^{2},\mathbf{f}_{i}^{3}\right],\mathbf{p}_{i}^{*}\right)$ denotes the primitive feature interpolated at the point $\mathbf{p}_{i}^{*}$ of maximum response, $\mathcal{G}$ is the set of primitives intersected by the ray, and $\alpha_{i}$ and $T_{i}$ correspond to the primitive opacity and accumulated transmittance defined in Eq.˜2.
5 Results
We present our results for radiance-field novel-view synthesis. Additionally, we demonstrate our method on two other reconstruction tasks: semantic scene reconstruction and high-resolution image fitting.
5.1 Radiance Field Reconstruction
5.1.1 Optimization
We directly adopt the densification strategy, loss function and regularization terms of 3DGS-MCMC [3dgsmcmc]. Specifically, our loss function is a combination of $\mathrm{L}_{1}$, D-SSIM, and regularizers on primitive opacity and scale:
$\mathcal{R}_{\alpha}=\frac{1}{P}\sum_{i=1}^{P}\alpha_{i}\,,\qquad\mathcal{R}_{s}=\frac{1}{P}\sum_{i=1}^{P}\lVert\mathbf{s}_{i}\rVert_{1}\,$, (4)
where $P$ is the number of primitives and $\mathbf{s}_{i}$ is the scale vector of primitive ii. Our final loss function is:
$\mathcal{L}=(1-\lambda)\,\mathcal{L}_{\mathrm{L}_{1}}+\lambda\,\mathcal{L}_{\text{D-SSIM}}+\lambda_{\alpha}\,\mathcal{R}_{\alpha}+\lambda_{s}\,\mathcal{R}_{s}\,.$ (5)
Similarly to the exponential scheduling applied to primitive positions [kerbl20233Dgaussians], we use a cosine annealing scheduler for the feature and MLP learning rates.
We also apply an Exponential Moving Average (EMA) filter [mueller2022instant] to the MLP weights $\theta$. This enhances robustness to noise and avoids overfitting to individual frames:
$\bar{\theta}_{t}\leftarrow\gamma\,\bar{\theta}_{t-1}+(1-\gamma)\,\theta_{t}\,$, (6)
where $\bar{\theta}_{t}$ denotes the filtered weights at step $t$ and $\gamma$ is the decay factor.
Finally, we dedicate the last 3000 training iterations to optimize only the feature and MLP weights, disabling all regularization terms and freezing all other parameters. We found that this refinement step slightly improves color fidelity, particularly in large-scale scenes. All hyperparameters and other smaller details will be included in the supplementary.
5.1.2 Implementation details
We implement our method in gsplat [ye2025gsplat], using the 3DGUT [3dgut] formulation, to which we add custom CUDA kernels for the forward and backward logic described in Sec.˜4. To reduce register pressure, we use half-precision memory fetches on feature vectors during rasterization, for both forward and backward kernels. Our neural network uses tiny-cuda-$n$’s [mueller2022instant] JIT-compiled cooperative vector MLPs, trained and evaluated in half precision (FP16). We apply an automatic scaling factor to the loss values in order to improve half-precision training stability [micikevicius2017mixed].
5.1.3 Results
We present the results of our method against a number of previous works using pure neural fields [barron2022mipnerf360], hashgrid-encoded neural fields [mueller2022instant, zipnerf], pure primitive-based methods [2dgs, 3dgsmcmc, 3dgut, betasplatting, trianglesplatting] with regular Spherical Harmonics (SH) appearance or more complex appearance models [sphericalVoronoi, texturedgaussians], and closer to us, methods mixing primitives (e.g. for acceleration) and neural fields [nestsplatting, radiancemesh] (Tab.˜1). In this table, we use the originally provided JPEG images for proper comparison with previous works in MipNeRF360; every other result presented in this section, unless otherwise stated, uses gsplat’s default downscaling for all methods tested. We measure the performance delta of this choice in the supplementary (Tab.˜10). We consistently outperform all previous works in a varied set of real datasets [barron2022mipnerf360, Knapitsch2017, hedman2018deep]. We include a selection of views in Fig.˜7. Our method particularly excels at high-frequency detail modelling, capturing specular highlights and reflections with superior fidelity.
Table 1: Comparison on MipNeRF360 [barron2022mipnerf360], Tanks & Temples [Knapitsch2017], and Deep Blending [hedman2018deep], on Neural Field-style methods, primitive-based methods, and mixed neural field/primitive-based methods. For the MipNeRF360 datasets, we disable gsplat’s default downscaling and instead train and evaluate directly on the provided JPEG-compressed reference images following prior work. We measure the effect in Tab.˜10. For our method, we use 64 features per primitive (16 per vertex), with a 128-wide $\times$ 3 hidden layers MLP, which results in a roughly similar number of total parameters as previous primitive-based approaches on average. For indoor scenes, which typically span a smaller spatial volume, we use 2M primitives, while for larger outdoor scenes we increase the primitive count to 5M. To regularize training in high-parallax regimes (e.g., indoor datasets), we encode the central camera ray rather than the individual ray direction. Finally, to ensure similar training epochs, we train indoor datasets longer (45k iterations for $\sim 290$ images) than outdoor (25k for $\sim 175$ images), following Spherical Voronoi [sphericalVoronoi]. Unlike Spherical Voronoi, we do not fine tune learning rates per dataset. Note that convergence of an MLP-based appearance model can be slower than SH-based ones.
MipNeRF360 Tanks & Temples Deep Blending
[…]
Instant NGP-Big [mueller2022instant] 25.59 0.695 0.375 21.92 0.740 0.342 24.96 0.815 0.459
Mip-NeRF 360 [barron2022mipnerf360] 27.60 0.788 0.275 22.22 0.754 0.290 29.40 0.899 0.306
ZipNeRF [zipnerf] 28.55 0.829 0.218 23.64 0.836 0.179
2DGS [2dgs] 27.22 0.804 0.275 22.85 0.827 0.244 29.56 0.904 0.325
3DGS-MCMC [3dgsmcmc] 27.99 0.830 0.229 24.46 0.866 0.174 29.49 0.912 0.306
3DGUT-MCMC [3dgsmcmc] 27.82 0.826 0.233 24.20 0.861 0.180 29.87 0.913 0.309
Beta Splatting-MCMC [betasplatting] 28.12 0.831 0.238 24.54 0.866 0.196 29.56 0.907 0.316
Spherical Voronoi [sphericalVoronoi] 28.56 0.835 0.228 24.80 0.871 0.172 30.34 0.914 0.299
Triangle Splatting [trianglesplatting] 27.00 0.808 0.231 23.05 0.843 0.191 28.92 0.891 0.308
Textured Gaussians [texturedgaussians] 27.35 0.827 24.26 0.854 28.33 0.891
NeST [nestsplatting] 26.54 0.776 0.260
Radiance Meshes [radiancemesh] 27.15 0.810 0.274 23.13 0.851 0.200 29.39 0.901 0.362
Neural Harmonic Textures (Ours) 28.74 0.834 0.216 25.68 0.882 0.141 30.94 0.919 0.302
In order to completely isolate the impact of our method, we include a controlled experiment in Table 2. We compare between 3DGS-MCMC, 3DGUT-MCMC, and different appearance models for the latter: regular SH, current state-of-the-art Spherical Voronoi functions [sphericalVoronoi] (SV) and ours (NHT). We implement all 4 under the same framework (gsplat), under strictly controlled conditions to ensure the only difference between them is the choice of appearance model itself. Table 2 shows that our method consistently outperforms both SH and SV across all benchmarks, without sacrificing real-time performance.
[…]
MipNeRF360 Tanks & Temples Deep Blending
[…]
3DGS + SH 27.94 0.829 0.246 251 24.25 0.861 0.188 294 29.98 0.912 0.317 331
3DGUT + SH 27.93 0.828 0.247 201 23.99 0.859 0.192 245 30.21 0.913 0.318 282
3DGUT + SV 28.15 0.823 0.248 202 24.18 0.861 0.187 242 30.29 0.912 0.320 267
3DGUT + NHT (Ours) 28.46 0.830 0.232 140 24.79 0.875 0.169 226 30.88 0.918 0.311 240
[…]
Figure 7: Comparison between our and previous works on radiance field reconstruction on scenes from MipNeRF360 [barron2022mipnerf360], Tanks and Temples [Knapitsch2017]. Our method models high frequency detail and view dependent effects to a higher degree than previous works.
Per-scene breakdowns and more details are provided in the appendix (Appendix Tab.˜11 and Tab.˜12).
5.1.4 Ablations
We ablate the key design decisions of our method on the MipNeRF360 dataset.
Optimization choices.
Appendix Tab.˜15 isolates the individual contributions of the training strategies detailed in Sec.˜5.1.2: learning rate scheduling, EMA on the MLP weights, loss scaling, the color-refinement phase, and opacity and scale regularization.
Primitive count.
Fig.˜6 illustrates how reconstruction quality scales with the number of primitives, ranging from 1K to 4M. Our method improves over previous works across this entire range, excelling particularly in highly compact regimes with low primitive counts. For instance, we achieve performance comparable to previous methods using 1M primitives with only a third of that amount, offering more flexibility in the memory-speed-quality trade-off.
Tab.˜20 and Tab.˜20 in the appendix explore the impact of varying the per-primitive feature dimension and the MLP architecture, respectively, to illustrate the resulting quality-speed trade-offs.
Feature encoding.
Tab.˜5 compares different encoding functions applied to the interpolated features.
Table 5: Ablating the choice of feature encoding function (MipNeRF360 dataset).
[…]
None (identity) 28.21 0.819 0.255
ReLU 28.19 0.819 0.255
Cosine 28.34 0.826 0.234
Cosine & Sine 28.46 0.830 0.232
5.2 Other Applications
5.2.1 Alternative primitive-based methods
We demonstrate the generality of our approach by incorporating Neural Harmonic Textures into 2DGS [2dgs] (as implemented in gsplat) and Triangle Splatting [trianglesplatting]. Our method is readily adapted to these 2D primitives by simply replacing the bounding tetrahedra with virtual triangles (2DGS) or directly anchoring feature vectors to explicit triangle geometry (Triangle Splatting), therefore using three feature vectors per primitive instead of four. The results are shown in Fig.˜6, with further details in the Supplementary.
5.2.2 Semantic Field Reconstruction
Our method can easily model higher dimensional data as well. By simply expanding the decoding head of our MLP, we can fit signals of arbitrary dimension. As the rest of the MLP layers remain small, and the size of feature vectors remains the same, the model is forced to exploit potential correlations between dimensions.
A particularly challenging case is semantic scene reconstruction, where fields are upwards of 512-dimensional. 3DGS-based methods relying on high-dimensional explicit features generally struggle with such scale, even when paired with feature upsampling networks.
We thus test our method on joint novel-view synthesis of RGB radiance and LSEG [LSEG] 512-wide semantic features. Due to the limited resolution of LSEG feature maps, we render and predict at the RGB resolution, but bilinearly downsample the spatial dimension of the resulting feature maps in order to supervise on ground truth features. We compare against Feature 3DGS [feature3dgs], a hybrid method relying on 128-wide semantic feature vector per primitive plus a CNN to jointly downsample the spatial dimension and upsample the feature dimension of the resulting rasterized image up to the full 512 dimensions, and another set of 48 SH basis weights for regular RGB color. Results for the aggregated MipNeRF360 dataset can be seen in Tab.˜6. We afford Feature-3DGS twice our training time budget, up to 7K iterations, slightly more than in the original paper (5K), to accommodate for the larger-scale datasets. We further ablate the per-primitive feature dimensionality for this task in Appendix Tab.˜13. Our results show substantial improvement over Feature 3DGS, and the capability to easily scale upwards or downwards in performance or memory, in a completely detached manner from the nature of the data being encoded. It is particularly interesting to note how small the drop in RGB quality is, which points to either high correlation among feature vectors or excessive free bandwidth in our RGB experiments. Our features can then be used in downstream tasks such as scene editing or semantic segmentation.
[…]
Feature 3DGS [feature3dgs] 26.27 0.785 0.262 43.68 0.985 3.02
Neural Harmonic Textures (Ours) 28.16 0.824 0.228 46.90 0.993 28.11
5.2.3 2D Image Reconstruction
Finally, we show results on 2D image fitting. To adapt our method to this application, we make two key changes with respect to the primitive-based approaches. First, we employ a connected 2D triangle mesh topology. This mesh is fully opaque: there is no alpha blending, no Gaussian or window kernel, simply interpolation of local features for a given pixel. Second, we change our barycentric interpolation to Clough-Tocher cubic interpolation. This makes interpolated features continuous over the triangle edges, which removes high frequency artifacts. The rest of the architecture is the same as described in Sec.˜4, and we include further details on the optimization strategy in the Supplementary.
We evaluate image reconstruction performance on a new curated dataset of high-resolution (45.7 MP) 14-bit HDR RAW images, and compare it against Instant NGP [mueller2022instant] and the JPEG-XL encoder [jpegxl], which supports high-dynamic range. We measure reconstruction quality at two compression ratios ($10\times$ and $100\times$), computing PSNR in $\mu$-law transformed space (PSNRμ), standard in HDR imaging pipelines. We also report PSNR, SSIM, and LPIPS on tonemapped images (subscript “tm”).
Results are shown in Tab.˜7. Our method achieves comparable pixel-error quality to Instant NGP in both HDR and linear spaces, while substantially reducing perceptual error: at $100\times$ compression, NHT achieves an LPIPS of 0.048 compared to 0.078 for Instant NGP, a $38\%$ relative improvement. JPEG-XL however still comes out on top, particularly in perceptual metrics, which is to be expected given its human-perception-based inductive biases. We can further compress our representation by a factor of 3 without noticeable loss in quality by aggressively reducing floating point precision of the implicit features, among other strategies. We include more details on this in the supplementary.
Table 7: High-resolution (45.7 MP) 14-bit HDR RAW image fitting, averaged over 15 images. We report PSNR in $\mu$-law transformed space (PSNRμ), common in HDR pipelines. PSNR, SSIM, and LPIPS are also computed on tonemapped images (subscript “tm”). All methods use the same memory budget per compression ratio.
[…]
$10\times$ JPEG-XL [jpegxl] 45.31 44.66 0.994 0.001
Instant NGP [mueller2022instant] 37.63 38.91 0.949 0.061
NHT (Ours) 37.16 39.87 0.963 0.023
$100\times$ JPEG-XL [jpegxl] 38.53 35.54 0.971 0.017
Instant NGP [mueller2022instant] 35.17 36.39 0.923 0.078
NHT (Ours) 35.06 36.43 0.927 0.048
[…]
6 Discussion
Limitations and future work
Our method focuses on increasing per-primitive expressivity, which has inherent downsides. In particular, our method tends to overfit more to individual views in scenarios with very sparse supervision, leading to lower novel view synthesis performance. Furthermore, while our method consistently achieves 140+ frames per second at inference time, the additional neural decoder results in slightly slower rendering than pure 3DGS methods(Tab.˜2).
In future work, we would like to investigate the feasibility of extracting automatic Level of Detail from a finer-grained harmonic decomposition. Given the generality of Neural Harmonic Textures, applications beyond novel view synthesis such as radiance caching, neural physically-based rendering and geometric reconstruction are also of great interest. Finally, our work opens the possibility of removing kernel functions entirely, as we can still propagate derivatives to explicit geometry without the spatially decaying opacity function, potentially enabling drastic performance gains.
Conclusion
We presented Neural Harmonic Textures, which combine particle-anchored feature vectors, harmonic activations, and a single image-space decoding pass to achieve state-of-the-art results in novel view synthesis. Our method consistently outperforms baselines across different particle counts, trains rapidly, and has high inference performance.
Moreover, it is compatible with existing deferred-shading rendering pipelines and supports a wide range of applications, from lifting semantic fields to high-resolution image reconstruction and beyond.
Acknowledgements
We would like to thank Ruilong Li and Martin Bisson for helpful discussions. Jorge Condor and Piotr Didyk acknowledge funding from the Swiss National Science Foundation (SNSF, Grant 200502) and an academic gift from Meta.
Supplementary Material
Appendix 0.A Additional Results
In Tab.˜11 we provide a detailed breakdown of Tab.˜2 from the main paper, with per-scene results on MipNeRF360, Tanks and Temples and Deep Blending. For this comparison, we isolate completely the effect of our method by ensuring an even setup for all methods. All four approaches are implemented under the same framework (gsplat), trained for 30k iterations, using 1M primitives for all scenes, sharing training parameters among all scenes, and limiting to 48 features per primitive reserved for appearance. All results presented in this paper compute the LPIPs metric using the VGG model with normalized inputs to the range of [1,1][-1,1]. Furthermore, in Tab.˜8 we provide another standardized comparison on all three benchmarks, only this time we match the same primitive count in each scene as the high-quality reported baselines in their respective papers, training for the same amount of iterations. We include our method with both 48 and 64 features per primitive (48F, 64F).
Table 8: Standardized comparison on MipNeRF360 [barron2022mipnerf360], Tanks & Temples [Knapitsch2017], and Deep Blending [hedman2018deep]: 3DGS-MCMC, 3DGUT-MCMC, and ours with 48 and 64 features per primitive (48F, 64F). All methods use the same primitive count (matching the high-quality preset in 3DGS-MCMC, which replicates the number of primitives in each scene from original 3DGS), 30k iterations, and the same hyperparameters across all scenes. Extended version of Tab.˜2 of the main paper. While our work particularly excels at extracting higher quality from less primitives, we still outperform previous works at high primitive counts across the board.
MipNeRF360 Tanks & Temples Deep Blending
[…]
3DGS-MCMC + SH 28.21 0.841 0.214 24.46 0.866 0.174 29.49 0.912 0.306
3DGUT-MCMC + SH 28.08 0.837 0.218 24.20 0.861 0.180 29.87 0.913 0.309
3DGUT-MCMC + NHT (Ours, 48F) 28.65 0.834 0.212 25.41 0.878 0.156 30.67 0.918 0.302
3DGUT-MCMC + NHT (Ours, 64F) 28.68 0.832 0.211 25.64 0.879 0.154 30.69 0.918 0.301
Table 9: Per-scene results for the standardized comparison (Tab.˜8). All methods use the same primitive count (matching the baseline default), 30k iterations, and the same hyperparameters across all scenes.
3DGS-MCMC + SH 3DGUT-MCMC + SH NHT (Ours, 48F) NHT (Ours, 64F)
[…]
MipNeRF360 – Outdoor
bicycle 25.87 0.801 0.183 25.69 0.790 0.193 25.53 0.779 0.202 25.62 0.776 0.202
garden 28.11 0.880 0.107 27.84 0.874 0.111 28.33 0.878 0.102 28.23 0.876 0.103
stump 27.32 0.808 0.203 27.15 0.802 0.215 27.06 0.788 0.215 27.02 0.786 0.215
treehill 23.29 0.670 0.310 23.13 0.664 0.313 21.96 0.639 0.311 22.27 0.640 0.308
flowers 22.25 0.652 0.312 22.25 0.648 0.318 21.90 0.637 0.315 21.77 0.631 0.312
MipNeRF360 – Indoor
room 32.56 0.939 0.240 32.43 0.939 0.241 33.82 0.945 0.229 33.75 0.946 0.227
counter 29.53 0.925 0.219 29.55 0.925 0.220 30.80 0.931 0.205 30.92 0.931 0.203
kitchen 32.10 0.938 0.136 31.96 0.938 0.136 33.43 0.944 0.127 33.44 0.944 0.124
bonsai 32.84 0.955 0.216 32.76 0.955 0.215 34.99 0.963 0.202 35.07 0.962 0.201
Tanks & Temples
train 22.48 0.832 0.221 22.19 0.826 0.229 23.88 0.856 0.198 24.33 0.858 0.194
truck 26.43 0.899 0.126 26.21 0.896 0.131 26.94 0.901 0.115 26.96 0.901 0.114
Deep Blending
drjohnson 29.25 0.910 0.310 29.13 0.910 0.315 30.06 0.916 0.308 30.09 0.916 0.306
playroom 29.72 0.913 0.301 30.60 0.916 0.302 31.28 0.920 0.297 31.28 0.920 0.296
M360 Outdoor Avg 25.37 0.762 0.223 25.21 0.756 0.230 24.96 0.744 0.229 24.98 0.742 0.228
M360 Indoor Avg 31.76 0.939 0.203 31.67 0.939 0.203 33.26 0.946 0.191 33.30 0.946 0.189
M360 Total Avg 28.21 0.841 0.214 28.08 0.837 0.218 28.65 0.834 0.212 28.68 0.832 0.211
T&T Avg 24.46 0.866 0.174 24.20 0.861 0.180 25.41 0.879 0.157 25.65 0.879 0.154
DB Avg 29.49 0.912 0.306 29.87 0.913 0.309 30.67 0.918 0.302 30.69 0.918 0.301
Tab.˜12 details the per-scene results of main Tab.˜1, which uses our split indoor/outdoor training configuration. This is our higher quality setup, which adapts to the different scene scale and dataset sizes of outdoor and indoor datasets.
Effect of reference image format.
As noted in Tab.˜1, MipNeRF360 images are distributed as pre-downscaled, re-compressed JPEG files. The default gsplat pipeline downscales and converts these to lossless PNGs before training and evaluation, which removes JPEG compression artifacts from the ground truth. In Tab.˜10 we compare NHT results under both protocols. Training and evaluating directly on the original JPEG references yields slightly lower PSNR (-0.28 dB), SSIM (-0.011), and higher LPIPS (++0.014) on average, as JPEG compression artifacts in the ground truth lower the effective quality ceiling.
Table 10: Effect of reference image format on MipNeRF360 evaluation for NHT (Ours). Downscaled PNGs uses gsplat’s default strategy, which downscales and converts the provided JPEG references to lossless PNGs before training and evaluation. JPEG (original) trains and evaluates directly on the provided JPEG-compressed references without any preprocessing. JPEG compression artifacts in the ground truth lower the effective quality ceiling, slightly degrading all three metrics.
[…]
Downscaled PNGs 29.022 0.8447 0.2026
JPEG (original) 28.738 0.8341 0.2164
$\Delta$ -0.284 -0.0106 ++0.0138
Table 11: Per-scene results on MipNeRF360 [barron2022mipnerf360], Tanks & Temples [Knapitsch2017], and Deep Blending [hedman2018deep], from Tab.˜2. This is the standardized setup limited to 1M primitives, 30k iterations, same number of appearance features per primitive (48) and same hyperparameters for all scenes, isolating entirely the effect of our approach from other variables.
3DGS-MCMC 3DGUT-MCMC + SH 3DGUT-MCMC + SV 3DGUT-MCMC + NHT (Ours)
[…]
MipNeRF360 – Outdoor
bicycle 25.61 0.774 0.246 25.51 0.771 0.250 25.21 0.758 0.255 25.43 0.771 0.243
garden 27.29 0.855 0.155 27.21 0.854 0.156 27.48 0.856 0.156 28.04 0.872 0.122
stump 26.93 0.788 0.250 26.93 0.789 0.251 25.95 0.757 0.270 26.58 0.772 0.249
treehill 23.45 0.665 0.366 23.30 0.663 0.368 23.16 0.653 0.371 23.13 0.660 0.340
flowers 22.04 0.628 0.362 22.07 0.627 0.363 21.76 0.616 0.367 21.60 0.615 0.351
MipNeRF360 – Indoor
room 32.33 0.938 0.245 32.36 0.937 0.246 32.67 0.939 0.243 33.03 0.943 0.234
counter 29.48 0.924 0.223 29.47 0.923 0.224 30.69 0.932 0.213 30.83 0.931 0.208
kitchen 31.67 0.935 0.144 31.72 0.935 0.144 32.38 0.938 0.140 32.98 0.942 0.132
bonsai 32.69 0.953 0.220 32.76 0.953 0.220 34.07 0.959 0.215 34.55 0.961 0.208
Tanks & Temples
train 22.50 0.832 0.224 22.10 0.827 0.231 22.32 0.831 0.223 23.11 0.851 0.202
truck 26.00 0.891 0.151 25.88 0.890 0.153 26.04 0.891 0.151 26.48 0.898 0.135
Deep Blending
drjohnson 29.50 0.909 0.320 29.60 0.909 0.321 29.69 0.909 0.321 30.24 0.915 0.317
playroom 30.45 0.915 0.313 30.82 0.916 0.315 30.89 0.915 0.318 31.52 0.920 0.306
M360 Outdoor Avg 25.06 0.742 0.276 25.00 0.741 0.278 24.71 0.728 0.284 24.96 0.738 0.261
M360 Indoor Avg 31.54 0.938 0.208 31.58 0.937 0.209 32.45 0.942 0.203 32.85 0.944 0.196
M360 Total Avg 27.94 0.829 0.246 27.93 0.828 0.247 28.15 0.823 0.248 28.46 0.830 0.232
T&T Avg 24.25 0.861 0.188 23.99 0.859 0.192 24.18 0.861 0.187 24.79 0.875 0.169
DB Avg 29.98 0.912 0.317 30.21 0.913 0.318 30.29 0.912 0.320 30.88 0.918 0.311
Table 12: Per-scene results of NHT (Ours) with split training configurations on MipNeRF360 [barron2022mipnerf360], Tanks & Temples [Knapitsch2017], and Deep Blending [hedman2018deep]. We adapt primitive count and training length per dataset to accommodate different dataset sizes and spatial scales, but keep the same learning rates and other hyperparameters. Indoor datasets benefit from encoding the central camera ray, instead of per-ray directions, due to their high parallax, acting as a regularization on ray orientation. Trained and evaluated on the original JPEG-compressed reference images.
[…]
M360 Outdoor (5M, 25k iters, per-ray)
garden 28.06 0.875 0.108
bicycle 25.53 0.782 0.203
stump 27.10 0.797 0.206
treehill 23.09 0.652 0.298
flowers 22.08 0.642 0.304
M360 Indoor (2M, 45k iters, center ray)
bonsai 35.09 0.957 0.222
counter 30.83 0.925 0.220
kitchen 33.40 0.940 0.138
room 33.43 0.936 0.250
Tanks & Temples (2.5M, 40k iters, center ray)
truck 26.91 0.900 0.112
train 24.45 0.865 0.169
Deep Blending (2M, 30k iters, center ray)
drjohnson 30.43 0.918 0.309
playroom 31.45 0.921 0.296
M360 Outdoor Avg 25.17 0.749 0.224
M360 Indoor Avg 33.19 0.940 0.207
M360 Total 28.74 0.834 0.216
T&T Avg 25.68 0.882 0.141
DB Avg 30.94 0.919 0.302
0.A.1 Semantic Reconstruction
Semantic Reconstruction Details
For our semantic reconstruction task, following a similar setup to Feature 3DGS [feature3dgs], we train a joint RGB and LSEG semantic feature field. Joint RGB and LSEG training uses the same approach, densification, and regularization as standard radiance-field training, with the differences being in the addition of semantic supervision losses and an expanded decoding head in our MLP. The deferred MLP is extended to predict both RGB (3 channels) and LSEG features (512 dimensions) via addition of an extra Pytorch linear layer. Rendering speed becomes bottlenecked by this linear layer, and more efficient fused alternatives or upsampling strategies could be explored in the future—we consider this a proof-of-concept implementation.
The total loss adds an LSEG term to the usual $\mathrm{L}_{1}$, D-SSIM, opacity and scale regularizers terms:
$\mathcal{L}_{\text{LSEG}}=\lambda_{\text{LSEG}}\,\mathcal{L}_{\mathrm{L}_{1}}\left(\hat{f},f^{*}\right)+\lambda_{\text{cos}}\,\left(1-\cos(\hat{f},f^{*})\right),$ (7)
where $\hat{f}$ and $f^{*}$ are predicted and ground-truth LSEG feature maps, and coscos their cosine similarity. We compute these losses at the native LSEG resolution: the model renders at RGB resolution (high); then, we bilinearly downsample the predicted feature map to match the (lower-resolution) LSEG targets. Tab.˜14 lists the LSEG-specific hyperparameters used in our final training. We largely kept the same configuration from our pure RGB radiance field experiments; further tuning for this specific use case may render higher performance.
Differences to RGB-only training.
Aside from the extra output head and loss terms, training proceeds as in the main paper: same optimizer, learning-rate schedules, EMA on the MLP, and color-refinement phase. The only architectural change is the increased output dimension (3 + 512 when fused, or a separate 512-dim linear layer). We do not use a separate training stage for LSEG; RGB and LSEG are optimized jointly from the start.
PCA visualization.
LSEG feature maps are 512-dimensional and therefore not directly viewable. For qualitative inspection, we project both ground-truth and predicted features to RGB via PCA. The PCA basis is computed from the first image, then applied the to all frames. Fig.˜9 shows example PCA visualizations. They allow a direct visual comparison of semantic structure and show that the predicted features align well with the target, while being substantially faster to generate and higher resolution than inferring the RGB frame only and running the LSEG model to obtain its feature map. Semantic scene reconstruction extracts detail from the multiple training views, improving the granularity of the semantic information available for downstream applications.
Finally, in Tab.˜13 we ablate the feature dimensionality for the joint RGB and LSEG semantic reconstruction task.
GT Ours GT Ours
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 9: LSEG feature PCA visualizations on test views from MipNeRF360. Rows from top to bottom: bicycle, garden, bonsai, kitchen. Each pair shows the ground-truth LSEG features (GT) alongside our rendered features (Ours). Our method faithfully reconstructs semantic feature maps while preserving sharp boundaries, at higher resolutions, and real-time.
[…]
4 23.81 0.853 0.254 44.33 0.988 76
8 25.40 0.876 0.200 47.10 0.994 75
16 26.06 0.885 0.178 48.23 0.995 72
32 26.34 0.887 0.166 48.68 0.996 72
64 26.42 0.887 0.163 48.79 0.996 70
Table 14: LSEG-specific hyperparameters used for joint RGB and semantic feature training.
Parameter Value
LSEG output feature dimension 512
LSEG $\mathrm{L}_{1}$ loss weight $\lambda_{\text{LSEG}}$ 1.0
LSEG cosine similarity weight $\lambda_{\text{cos}}$ 0.1
Per-primitive feature dim. (shared with RGB) 80
Appendix 0.B Feature interpolation
For a spatial coordinate $\mathbf{p}$ lying on the surface of our topology, or inside of its volume, we can compute the feature vector by barycentrically interpolating the feature vectors at the vertices of the shape. For example, for a 2D triangle with vertices $\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{v}_{3}$, the feature vector $f(\mathbf{p})$ is computed as
$f(\mathbf{p})=\sum_{i=1}^{n}\alpha_{i}f_{i}$ (8)
where $f_{i}$ are the features at the triangle vertices, and $\alpha_{i}$ are the barycentric coordinates at the evaluated point. In the case of 2D triangles, barycentric weights are computed as
$\alpha_{i}=\frac{A_{i}}{A},\quad A=\sum_{j=1}^{n}A_{j}$ (9)
where AA is the total area of the triangle and $A_{i}$ is the area of the sub-triangle formed by the evaluation point $\mathbf{p}$ and the edge opposite to vertex ii. For a triangle with vertices $\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{v}_{3}$:
$A_{1}$ $=\frac{1}{2}\left\|(\mathbf{v}_{2}-\mathbf{p})\times(\mathbf{v}_{3}-\mathbf{p})\right\|,$
$A_{2}$ $=\frac{1}{2}\left\|(\mathbf{v}_{3}-\mathbf{p})\times(\mathbf{v}_{1}-\mathbf{p})\right\|,$
$A_{3}$ $=\frac{1}{2}\left\|(\mathbf{v}_{1}-\mathbf{p})\times(\mathbf{v}_{2}-\mathbf{p})\right\|.$ (10)
For 2D connected triangles, this is similar to vertex features recently introduced in the context of neural global illumination [ruisu2025]. Our formulation however is more general and extends to higher order topologies and disconnected meshes. For tetrahedra, barycentric weights are similarly computed as
$\alpha_{i}=\frac{V_{i}}{V},\quad V=\sum_{j=1}^{n}V_{j}$ (11)
where $V$ is the total volume of the tetrahedron and $V_{i}$ is the volume of the sub-tetrahedron formed by the evaluation point $\mathbf{p}$ and the face opposite to vertex ii. For a tetrahedron with vertices $\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{v}_{3},\mathbf{v}_{4}$:
$V_{1}$ $=\frac{1}{6}\left\|(\mathbf{v}_{2}-\mathbf{p})\times(\mathbf{v}_{3}-\mathbf{p})\cdot(\mathbf{v}_{4}-\mathbf{p})\right\|,$
$V_{2}$ $=\frac{1}{6}\left\|(\mathbf{v}_{3}-\mathbf{p})\times(\mathbf{v}_{4}-\mathbf{p})\cdot(\mathbf{v}_{1}-\mathbf{p})\right\|,$
$V_{3}$ $=\frac{1}{6}\left\|(\mathbf{v}_{4}-\mathbf{p})\times(\mathbf{v}_{1}-\mathbf{p})\cdot(\mathbf{v}_{2}-\mathbf{p})\right\|,$
$V_{4}$ $=\frac{1}{6}\left\|(\mathbf{v}_{1}-\mathbf{p})\times(\mathbf{v}_{2}-\mathbf{p})\cdot(\mathbf{v}_{3}-\mathbf{p})\right\|.$ (12)
0.B.1 Optimization choices and hyperparameters
Tab.˜15 ablates the individual contributions of the training strategies described in the main paper (learning rate scheduling, EMA on the MLP weights, loss scaling, color-refinement phase, and opacity and scale regularization). Tab.˜16 lists the full set of hyperparameters and optimization details used for radiance field reconstruction.
[…]
Our Model 28.49 0.828 0.153
A) No LR Scheduling 28.37 0.829 0.147
B) No EMA on MLP Weights 28.46 0.828 0.152
D) No Color Refinement Phase 28.35 0.827 0.154
E) No Opacity Regularization 27.85 0.806 0.181
F) No Scale Regularization 28.50 0.828 0.153
G) No Direction Encoding 28.11 0.816 0.155
I) No Direction Scaling 28.35 0.822 0.158
Table 16: Hyperparameters and optimization details for radiance field reconstruction (MipNeRF360, Tanks & Temples, Deep Blending). Values follow our gsplat implementation [ye2025gsplat].
Training
Color refinement phase Last 3 000 steps (geometry frozen, no reg.)
Loss & regularization
SSIM weight $\lambda$ (D-SSIM vs $\mathrm{L}_{1}$) 0.1
Opacity regularization $\lambda_{\alpha}$ 0.02
Scale regularization $\lambda_{s}$ 0.005
Learning rates (initial)
Positions (means) $1.6\times 10^{-4}$
Scales $5\times 10^{-3}$
Opacities $5\times 10^{-2}$
Quaternions $1\times 10^{-3}$
Deferred features $1.5\times 10^{-2}$
Deferred MLP $6.8\times 10^{-4}$
LR schedule
Positions Exponential decay (final factor 0.01)
Features & MLP Cosine annealing (final factor 0.1)
EMA (deferred MLP)
Decay $\gamma$ 0.95
Start step 0
Appearance (NHT)
Feature dim. per primitive 16-64
MLP hidden dim. $\times$ layers $64-128\times 2-3$
View encoding Spherical harmonics (2nd degree)
[…]
Appendix 0.C 2D Image Reconstruction Details
We provide additional details on the high-resolution HDR image fitting experiment of Sec.˜5.2.3. We adapt Neural Harmonic Textures to this domain while preserving the core idea: learnable per-vertex features, interpolated per pixel and activated with periodic functions, and decoded by a shallow MLP. However, there are a number of differences to the 3D and semantic reconstruction tasks, which we detail below.
0.C.1 Representation
Triangle mesh.
Unlike the volumetric disconnected primitives used in the scene reconstruction tasks, the 2D image fitting experiment employs a connected Delaunay triangulation that tiles the image plane without overlap. Every pixel belongs to exactly one triangle; there is no alpha blending, no Gaussian kernel, and no depth ordering. Vertices can be freely learned and positioned, and carry individual learnable feature vectors. Given a query pixel, its containing triangle is found via a tile-accelerated rasterizer, and the features at the triangle’s three vertices are interpolated to produce a per-pixel feature vector.
Clough–Tocher $C^{1}$ interpolation.
Standard barycentric (linear) interpolation produces a $C^{0}$ feature field: while continuous, the feature gradient is discontinuous across triangle edges. This turns out to be problematic when meshes are fully opaque, as it creates high frequency artifacts. Despite being followed by a frequency encoding and a nonlinear MLP, these gradient discontinuities manifest as visible seam artifacts along every edge.
To eliminate this, we replace linear interpolation with Clough–Tocher cubic interpolation, which constructs a degree-3 Bézier patch on each triangle from vertex values and vertex gradients, guaranteeing $C^{1}$ continuity everywhere (and $C^{\infty}$ within each triangle). The 10 Bézier control points are:
$c_{300}$ $=f_{0},\quad c_{030}=f_{1},\quad c_{003}=f_{2},$
$c_{210}$ $=f_{0}+\tfrac{1}{3}\,\nabla f_{0}\cdot\mathbf{e}_{01},\quad c_{120}=f_{1}-\tfrac{1}{3}\,\nabla f_{1}\cdot\mathbf{e}_{01},$
$c_{021}$ $=f_{1}+\tfrac{1}{3}\,\nabla f_{1}\cdot\mathbf{e}_{12},\quad c_{012}=f_{2}-\tfrac{1}{3}\,\nabla f_{2}\cdot\mathbf{e}_{12},$
$c_{102}$ $=f_{2}+\tfrac{1}{3}\,\nabla f_{2}\cdot\mathbf{e}_{20},\quad c_{201}=f_{0}-\tfrac{1}{3}\,\nabla f_{0}\cdot\mathbf{e}_{20},$
$c_{111}$ $=\tfrac{1}{6}\bigl(c_{210}{+}c_{120}{+}c_{021}{+}c_{012}{+}c_{102}{+}c_{201}\bigr)$
$\quad-\tfrac{1}{6}\bigl(c_{300}{+}c_{030}{+}c_{003}\bigr),$ (13)
where $f_{i}$ and $\nabla f_{i}$ are the feature value and gradient at vertex $i$, and $\mathbf{e}_{ij}=\mathbf{v}_{j}-\mathbf{v}_{i}$.
Vertex gradient estimation.
The Clough–Tocher scheme requires per-vertex gradients. We estimate these via a per-vertex least-squares fit over the one-ring neighborhood: for each vertex $v$, we gather all neighbor vertices $\{v_{j}\}_{j\in\mathcal{N}(v)}$ and solve
$\nabla f_{v}=\mathbf{A}_{v}^{-1}\,\mathbf{b}_{v},\quad\mathbf{A}_{v}=\sum_{j}\Delta\mathbf{p}_{j}\,\Delta\mathbf{p}_{j}^{\top},\quad\mathbf{b}_{v}=\sum_{j}\Delta f_{j}\,\Delta\mathbf{p}_{j},$ (14)
where $\Delta\mathbf{p}_{j}=\mathbf{p}_{j}-\mathbf{p}_{v}$ and $\Delta f_{j}=f_{j}-f_{v}$. The $\times2$ system is solved analytically per vertex in a CUDA kernel. This operation is differentiable and computed at every forward pass, enabling gradients to flow through the vertex topology.
Feature encoding and MLP.
Feature encoding and MLP decoding are the same as in the main scene reconstruction task.
0.C.2 Training
Pixel sampling.
At each iteration, we sample a batch of pixels via stratified sampling. Ground-truth values at sub-pixel coordinates are obtained via bilinear interpolation on the target image, implemented as a CUDA texture lookup for efficiency.
Optimizers and schedules.
We use Adam [kingma2017adam] with separate parameter groups for positions ($\text{lr}=10^{-4}$), features ($\text{lr}=5\times 10^{-3}$), and MLP weights ($\text{lr}=5\times 10^{-5}$). An exponential decay schedule is applied after 20 000 iterations with a factor of $0.33$ every 10 000 steps, following Instant NGP [mueller2022instant].
Loss function.
We train with a pixel-wise MSE loss in $\mu$-law encoded space. We found this to produce better results than using the relative L2 loss suggested in the Instant NGP paper. Raw 14-bit sensor values $x$ are transformed as $f(x)=\log(1+\mu\,x/w)/\log(1+\mu)$ with $\mu=5000$ and $w=2^{14}-1=16{,}383$ (the white level of 14-bit data). This compresses the dynamic range before supervision, placing relatively more emphasis on shadow detail, which is standard in HDR imaging pipelines.
0.C.3 Coarse-to-Fine Densification
Training begins from a sparse mesh initialized via edge-aware sampling: we compute Sobel gradient magnitudes on a downsampled version of the target image and sample initial vertex positions proportionally to edge strength, with a small uniform floor for coverage. The result is a Delaunay triangulation of the sampled points, with boundary vertices ensuring full image coverage.
The mesh is progressively refined through densification. Every 500 iterations (starting at iteration 1 500 and ending at 15 000), we score each triangle by
$\text{score}_{t}=\overline{\text{DSSIM}}_{t}\cdot|\text{pixels}_{t}|^{\alpha},$ (15)
where $\overline{\text{DSSIM}}_{t}$ is the mean DSSIM (per-pixel structural dissimilarity) of triangle $t$ and $\alpha=0.75$ is the area-weighting exponent. The area weighting prevents repeated splitting of tiny, high-error triangles and distributes vertices more evenly in proportion to their spatial impact. Triangles with fewer than 3 pixels are excluded from scoring.
At each densification step, we insert centroid vertices for the top-scoring triangles (up to $0.35\times V$ new vertices, where $V$ is the current count), initialize their features by averaging the parent triangle’s vertex features, and re-run Delaunay triangulation on the full vertex set. We also detect when the mesh becomes degenerate (e.g. due to vertices becoming colinear) and re-mesh, although this rarely occurs in practice.
0.C.4 CUDA Rasterizer
All rasterization and interpolation operations are implemented as custom forward and backward CUDA kernels, supporting both full-image rendering and batched pixel-query modes. Training generally takes between 1-3 minutes in an RTXA6000 Ada.
0.C.5 Post-Training Compression
We also briefly explore the use of post-training compression techniques to further reduce storage requirements.
We compress vertex positions to UInt16, vertex features to uniform Int8 (with a per-channel scale and offset), and MLP weights to FP16. An entropy coder (Zstandard, level 19) is applied to the serialized payload. Combining these techniques, the resulting images can be compressed by an additional factor of approximately $3\times$ beyond the training result, with negligible quality loss. Tab.˜21 summarizes the results of the post-training compression techniques.
[…]
0.C.6 Per-Image Results
Tabs.˜22 and 23 provide per-image results for the 2D image fitting experiment at $10\times$ and $100\times$ compression ratios, respectively. We provide visual comparisons in Fig.˜8 and Fig.˜10. Our method achieves competitive PSNR in both $\mu$-law and tonemapped spaces while substantially outperforming Instant NGP in perceptual quality (LPIPS). The advantage is most pronounced at high compression ratios, where the non-overlapping mesh topology and $C^{1}$ feature interpolation avoid the ringing and block artifacts that plague hash-grid methods.
[…]
Image NHT NGP JXL NHT NGP JXL NHT NGP JXL NHT NGP JXL
DSC_1012 33.98 34.04 37.70 36.92 37.59 30.38 0.97 0.96 0.99 0.0090 0.0137 0.0004
DSC_1764 40.06 39.04 49.03 43.61 41.73 32.02 0.98 0.97 0.99 0.0401 0.1149 0.0012
DSC_1967 33.60 35.01 41.48 36.50 36.19 27.92 0.95 0.93 0.99 0.0108 0.0178 0.0004
DSC_2172 35.83 36.59 41.74 40.23 38.95 40.31 0.97 0.95 0.99 0.0182 0.0317 0.0008
DSC_2655 41.80 40.91 50.26 44.38 42.73 55.95 0.98 0.98 1.00 0.0292 0.0850 0.0008
DSC_3095 38.07 38.78 46.05 39.83 38.43 52.47 0.96 0.94 1.00 0.0184 0.0450 0.0009
DSC_3351 40.00 42.20 48.40 44.53 42.97 55.37 0.98 0.98 1.00 0.0046 0.0179 0.0002
DSC_4748 33.39 34.68 37.97 37.53 37.08 31.97 0.95 0.94 0.99 0.0218 0.0544 0.0010
DSC_5007 36.09 36.80 47.75 37.46 36.54 49.40 0.93 0.90 0.99 0.0250 0.0430 0.0009
DSC_6718 41.30 40.15 51.27 44.57 41.64 55.28 0.98 0.97 1.00 0.0464 0.1345 0.0010
DSC_6926 31.24 33.15 34.93 33.27 34.28 27.42 0.93 0.91 0.98 0.0246 0.0449 0.0022
DSC_9008 40.78 40.10 49.73 41.93 39.98 53.74 0.98 0.96 1.00 0.0111 0.0493 0.0006
DSC_1658 39.05 38.51 50.39 41.68 39.57 53.53 0.97 0.95 1.00 0.0471 0.1445 0.0018
DSC_4176 35.17 36.89 48.17 36.50 36.64 50.77 0.95 0.93 1.00 0.0215 0.0597 0.0006
DSC_7685 37.11 37.54 44.83 39.08 39.33 53.30 0.97 0.96 1.00 0.0184 0.0550 0.0004
Average 37.16 37.63 45.31 39.87 38.91 44.66 0.96 0.95 0.99 0.0231 0.0608 0.0009
[…]
Image NHT NGP JXL NHT NGP JXL NHT NGP JXL NHT NGP JXL
DSC_1012 31.49 31.65 32.58 33.34 33.51 30.18 0.93 0.93 0.98 0.0243 0.0255 0.0065
DSC_1764 37.41 37.22 41.89 39.95 39.70 31.89 0.96 0.96 0.98 0.0552 0.1341 0.0344
DSC_1967 30.65 31.05 34.25 32.41 32.61 27.66 0.89 0.88 0.96 0.0328 0.0341 0.0057
DSC_2172 32.96 33.31 36.09 35.59 35.90 38.59 0.92 0.92 0.97 0.0419 0.0436 0.0182
DSC_2655 40.30 40.16 43.56 41.74 41.36 26.05 0.98 0.98 0.99 0.0396 0.0942 0.0121
DSC_3095 36.52 35.93 39.03 36.74 35.82 43.61 0.93 0.92 0.98 0.0459 0.0663 0.0184
DSC_3351 40.79 41.23 43.53 42.12 42.51 29.01 0.98 0.98 0.99 0.0099 0.0193 0.0020
DSC_4748 31.05 31.26 32.83 34.23 34.38 31.47 0.91 0.91 0.96 0.0516 0.0724 0.0142
DSC_5007 33.88 34.00 38.84 33.65 33.68 36.35 0.85 0.84 0.95 0.0692 0.0729 0.0143
DSC_6718 38.55 38.68 43.99 40.74 40.19 47.59 0.96 0.96 0.99 0.0847 0.1571 0.0362
DSC_6926 29.48 29.72 30.01 29.91 30.54 27.02 0.87 0.86 0.93 0.0587 0.0629 0.0331
DSC_9008 38.38 38.39 43.15 37.90 37.92 34.54 0.95 0.95 0.98 0.0275 0.0514 0.0096
DSC_1658 36.56 36.45 41.84 37.87 37.29 44.60 0.93 0.93 0.98 0.0890 0.1768 0.0489
DSC_4176 33.49 33.56 38.28 33.60 33.63 40.03 0.90 0.89 0.97 0.0521 0.0880 0.0123
DSC_7685 34.42 34.91 38.10 36.63 36.86 44.49 0.94 0.94 0.98 0.0345 0.0681 0.0081
Average 35.06 35.17 38.53 36.43 36.39 35.54 0.93 0.92 0.97 0.0478 0.0778 0.0183
Instant NGP NHT (Ours) Reference
Refer to caption Refer to caption Refer to caption
42.5 dB / 0.019 42.1 dB / 0.010 DSC_3351
Refer to caption Refer to caption Refer to caption
40.2 dB / 0.157 40.7 dB / 0.085 DSC_6718
Refer to caption Refer to caption Refer to caption
39.7 dB / 0.134 39.9 dB / 0.055 DSC_1764
Refer to caption Refer to caption Refer to caption
30.5 dB / 0.063 29.9 dB / 0.059 DSC_6926
Refer to caption Refer to caption Refer to caption
36.9 dB / 0.068 36.6 dB / 0.035 DSC_7685
[…]
Tab.˜24 provides the full set of hyperparameters for the 2D image fitting experiment.
Table 24: Hyperparameters and optimization details for the 2D image fitting experiment. The parameter budget (maximum vertices, MLP size) is derived from the target compression ratio and image dimensions and bitrate.
Representation
Mesh topology Delaunay triangulation (shared vertices)
Interpolation Clough–Tocher $C^{1}$ cubic
Feature activation SinCos: $\sin(f),\cos(f)]$
Feature dim. per vertex 4-16
MLP hidden dim. 6412864-128 $\times$ 242-4 layers
Training
Total iterations 25 000
Pixel sampling Stratified, 160 000 pixels/batch
Loss function MSE
Learning rates (initial)
Vertex positions $1\times 10^{-4}$
Vertex features $5\times 10^{-3}$
MLP weights $5\times 10^{-5}$
LR schedule
All parameters Exponential decay:
$\eta(t)=\eta_{0}\cdot 0.33^{\max(0,\,(t-20\text{k})\,/\,10\text{k})}$
Densification
Strategy Coarse-to-fine Delaunay re-meshing
Schedule Steps 1 500–15 000, every 500 steps
[…]
Scoring DSSIM-based, area-weighted:
$\text{score}_{t}=\overline{\text{DSSIM}}_{t}\cdot|\text{pixels}_{t}|^{0.75}$
Min. triangle size 3 pixels (skip smaller triangles)
Initialization Edge-aware sampling (Sobel)
HDR pipeline
Input 14-bit RAW (Nikon NEF), 45.7 MP
Encoding space $\mu$-law: $f(x)=\log(1+\mu\,x/w)/\log(1+\mu)$
$\mu$ parameter 5 000
White level $w$ $2^{14}-1=16\,383$
GT pixel lookup Bilinear interpolation